OpenStudy (anonymous):

Which of the following represents inverse variation? y equals nine divided by x y equals three plus x y divided by x equals negative ten y equals negative one-sixth x

OpenStudy (tkhunny):

Which one includes division?

OpenStudy (anonymous):

Which value for slope represents the steepest line? Explain. Negative thirty-five eighths or negative three or negative five-sixths or negative four

OpenStudy (theeric):

Well, if there is "inverse variation" between \(x\) and \(y\), then \(x\times y\) equals a constant. http://www.mathwords.com/i/inverse_variation.htm

OpenStudy (anonymous):

ignore the last question help with this one

OpenStudy (anonymous):

???

OpenStudy (theeric):

These are your choices: \(\bullet\quad y=\dfrac{9}{x}\) \(\bullet\quad y=3+x\) \(\bullet\quad \dfrac{y}{x}=-10\) \(\bullet\quad y=-\dfrac{1}{6}x\) Now, I'm used to just looking for where \(y\) equals a term that has a plain old \(x\) in the denominator. Do you see an equation like that?